OpenStudy (anonymous):

Suppose a function does the following to any X you input into it 1.) Subtracts 3 2.) Multiplies by -1.3 3.) Takes the cube root Describe what the inverse function would do.

OpenStudy (ash2326):

inverse function will do the reverse a) add 3 b) divide by -1.3 c) will cube it

OpenStudy (anonymous):

Can you give an example using an equation?

OpenStudy (ash2326):

1) subtract 3 say y=x-3 find x from this x=y+3 replace x by y and y by x its inverse is y=x+3 2) y= x*(-1.3) y/(-1.3)=x replace x by y and y by x its inverse is y=x/(-1.3) 3)y= (x)^1/3 y^3=x replace x by y and y by x it's inverse is y=x^3

OpenStudy (ash2326):

did you get it?

OpenStudy (anonymous):

perfect thank you for clearing this up

OpenStudy (ash2326):

Glad to help :)